# New bounds for proper $h$-conflict-free colourings

### v1

@@ Abstract.
A proper $k$-colouring of a graph $G$ is called $h$-conflict-free if every vertex $v$ has at least $\min\,\{h,\deg(v)\}$ colours appearing exactly once in its neighbourhood. Let $\chi_{\rm pcf}^{h}(G)$ denote the minimum $k$ such that such a colouring exists. We show that for every fixed $h\geq 1$, every graph $G$ of maximum degree $\Delta$ satisfies $\chi_{\rm{pcf}}^{h}(G)\leq h\Delta+\mathcal{O}(\log\Delta)$. This expands on the work of Cho et al., and improves a recent result of Liu and Reed in the case $h=1$. We conjecture that for every $h\geq 1$ and every graph $G$ of maximum degree $\Delta$ sufficiently large, the bound $\chi_{\rm pcf}^{h}(G)\leq h\Delta+1$ should hold, which would be tight. When the minimum degree $\delta$ of $G$ is sufficiently large, namely $\delta\geq\max\{100h,3000\log\Delta\}$, we show that this upper bound can be further reduced to $\chi_{\rm{pcf}}^{h}(G)\leq\Delta+\mathcal{O}(\sqrt{h\Delta})$. This improves a recent bound from Kamyczura and Przybyło when $\delta\leq\sqrt{h\Delta}$.
@@ 1. Introduction
For a positive integer $k$, let us denote $[k]\coloneqq\{1,\dots,k\}$ the set of the first $k$ integers. A $k$-colouring of a graph $G$ is an assignment $\sigma\colon V(G)\rightarrow[k]$. A colouring $\sigma$ of $G$ is proper if $\sigma(u)\neq\sigma(v)$ for every edge $uv\in E(G)$. Given a colouring $\sigma$ and a vertex $v\in V(G)$, a witness of $v$ is a neighbour $u\in N(v)$ such that the colour $\sigma(u)$ appears exactly once in $N(v)$; we say that $\sigma(u)$ is a solitary colour of $v$ in $\sigma$. The colouring $\sigma$ is proper conflict-free (or simply pcf) if it is proper and every non-isolated vertex $v$ has a witness. We let $\chi_{\rm pcf}(G)$ be the smallest integer $k$ such that a pcf $k$-colouring of $G$ exists. The notion of pcf colouring was introduced by Fabrici et al. [8], and was first investigated in the class of graphs of maximum degree $\Delta$ by Caro, Petruševski and Škrekovski [3]. Among other results, they obtained the upper bound $\chi_{\rm pcf}(G)\leq 5\Delta(G)/2$, and conjectured that the optimal upper bound should actually be $\Delta(G)+1$.
@@
The condition $\Delta\geq 3$ is needed because the conjectured bound does not hold for cycles. Indeed $\chi_{\rm pcf}(C_{5})=5=\Delta(C_{5})+3$, and for every other cycle $C$ whose length is not a multiple of $3$, $\chi_{\rm pcf}(C)=4=\Delta(C)+2$. The condition that $G$ is connected ensures that there is no isolated vertex. The conjectured upper bound is best possible. For instance, if $G$ is a graph obtained by subdividing once any subset of edges from the complete graph $K_{n}$ with $n\geq 5$, then $\chi_{\rm pcf}(G)=n=\Delta(G)+1$. In particular, when $G={K_{n}}^{1/2}$, the difference between $\chi(G)$ and $\chi_{\rm pcf}(G)$ is unbounded; we have $\chi(G)=2$ while $\chi_{\rm pcf}(G)=n$.
As a first step towards resolving 1, Cranston and Liu [6] observed that a simple greedy algorithm provides a pcf $(2\Delta+1)$-colouring of any graph $G$ of maximum degree $\Delta$. They moreoever proved that $\chi_{\rm pcf}(G)<1.656\Delta$ when $\Delta$ is large enough. More recently, Liu and Reed [10] proved that 1 holds asymptotically, by showing $\chi_{\rm pcf}(G)\leq\Delta+\mathcal{O}(\Delta^{2/3}\log\Delta)$ as $\Delta\to\infty$. Our first contribution is to improve this upper bound by eliminating the polynomial factor in the second-order term.
@@
Fix any linear ordering $(v_{1},\dots,v_{n})$ of $V(G)$. For a vertex $v\in V(G)$, let $\mathcal{W}(v)$ be the first $\min\,\{h,\deg(v)\}$ neighbours of $v$ in the ordering. For $i$ from $1$ to $n$, colour $v_{i}$ with the first colour which does not appear in $N(v_{i})$, nor in $\mathcal{W}(u)$ for $u\in N(v_{i})$. Let $\sigma$ be the colouring obtained at the end of this procedure. One easily checks that $\sigma$ is proper. Furthermore, for all $v\in V$, the vertices of $\mathcal{W}(v)$ are witnesses of $v$ (the fact that $\mathcal{W}(v)$ is coloured before $N(v)\setminus\mathcal{W}(v)$ is important), therefore $\sigma$ is $h$-pcf. Observe that the number of colour constraints when colouring $v_{i}$ is at most $(h+1)\deg(v)$, therefore $\sigma$ uses at most $(h+1)\Delta+1$ colours. ∎
@@
Let $G$ be a graph of maximum degree $\Delta$, and let $21\leq h\leq\Delta-$ 2 be an integer. Then $\chi_{\rm pcf}^{h}(G)\leq(h+1)\Delta-1$.
For the specific case $h=\Delta-2$, they exhibit a graph $G$ of maximum degree $\Delta$ arbitrarily large such that $\chi_{\rm pcf}^{h}(G)=(h+1)(\Delta-1)$. Note that this last term is also equal to $h\Delta+1$. We consider the opposite end of the spectrum by considering $h$ a fixed constant, and make an asymptotic improvement on 4 in that case.
@@
$\chi_{\rm pcf}^{h}(G)\leq h\Delta+\mathcal{O}\left(\log\Delta\right).$
2 is a direct corollary of 5 with $h=1$. We note that the hidden constant in the term $\mathcal{O}(\log\Delta)$ has a linear dependency in $h$; this constant is explicitly stated in 24.
The notion of pcf colourings can also be seen as a strengthening of odd colourings, cf. [12, 4, 7]. For a graph $G$ and an integer $h\geq 1$, an $h$-odd colouring of $G$ is a proper colouring such that every vertex $v$ observes at least $\min\,\{h,\deg(v)\}$ colours in its neighbourhood that appear an odd number of times. Let $\chi_{o}^{h}(G)$ be the minimum number of colours required for an $h$-odd colouring of $G$, and $\chi_{o}(G)\coloneqq\chi_{o}^{1}(G)$. Clearly, $\chi_{o}^{h}(G)\leq\chi_{\rm pcf}^{h}(G)$ for every $h\geq 1$. A weaker version of 1 was first formulated in [2] for the odd chromatic number of graphs of maximum degree $\Delta$, and it was shown in [7, Theorem 3] that $\chi_{o}(G)\leq\Delta+\mathcal{O}\left(\log\Delta\right)$. In that sense, 2 is also a strengthening of this result.
If $h$ is a fixed integer and $G$ is a graph of maximum degree $\Delta$, then we have $\chi_{o}^{h}(G)\leq h\Delta+\mathcal{O}\left(\log\Delta\right)$ as a direct consequence of 5. The first-order term of this upper bound was shown to be tight in [7, Proposition 28].
@@
In turn, 6 demonstrates that 5 is tight up to the second-order term, i.e. for every fixed integer $h\geq 1$, there exists a graph of maximum degree $\Delta$ arbitrarily large such that $\chi_{\rm pcf}^{h}(G)\geq h\Delta+1$. In light of these observations, we propose an extended version of 1.
@@
The lower bound provided by 6 relies on a graph construction with many vertices of small degree (namely, of degree $h+1$). When there is an additional restriction on the minimum degree of $G$, it is possible to derive smaller upper bounds on $\chi_{\rm pcf}^{h}(G)$. In particular, Liu and Reed [11] have recently proved that 1 holds for regular graphs in a strong sense. Their result can be stated as follows.
@@
The condition on the minimum degree in 8 is of course very restrictive. One could wonder what happens when this condition is relaxed. This has been considered by Kamyczura and Przybyło [9], who showed the following.
@@
Let $G$ be a graph of maximum degree $\Delta$ and minimum degree $\delta$. For every integer $h$ satisfying $7520\log\Delta\leq h\leq\delta/$ 75, one has
@@
The condition on $h$ implies that $\delta>1500\log\Delta$. Note that the condition $h\geq 20\log\Delta$ in the above statement is non-restrictive, since $\chi_{\rm pcf}^{h}(G)$ is monotonous in $h$: when $h<20\log\Delta$, one can use $\chi_{\rm pcf}^{h}(G)\leq\chi_{\rm pcf}^{\left\lceil 20\log\Delta\right\rceil}(G)$ to obtain an upper bound on $\chi_{\rm pcf}^{h}(G)$ with an application of 9. A direct consequence of 9 is that, when $h$ is a fixed integer and $\delta\gg\log\Delta$, one has $\chi_{\rm pcf}^{h}(G)\leq\Delta+\mathcal{O}(\Delta\log\Delta/\delta)$. We improve this bound when the minimum degree is not too large.
@@
Let $G$ be a graph of maximum degree $\Delta$ and minimum degree $\delta$. For every integer $h$ satisfying $10030\log\Delta\leq h\leq\delta/$ 100, one has
@@
The upper bound provided by 10 is smaller than that of 9 in the regime $\delta=\mathcal{O}(\sqrt{h\Delta})$, and performs best at $\delta=100h$. However, we suspect that there is still room for improvement in that regime. For matters of comparison, it was proved in [7, Corollary 6.3] that if $23\log\Delta\leq h\leq\delta/$ 2, then $\chi_{o}^{h}(G)\leq\Delta+\mathcal{O}(h)$, and we conjecture that a similar bound should hold for $h$-pcf colourings as well.
@@
In some sense, our contribution in 10 is a halfway progression towards this conjecture, when $h=\Theta(\delta)$.
1.1. Notations and ouline
@@
The paper is organised as follows. In Section 2, we provide probabilistic tools that we will rely on in our proofs. We consider general graphs in Section 3 and prove 5. In Section 4 we consider graphs of constrained minimum degree and prove 10.
@@ 2. Probabilistic tools
In our proofs, we shall mostly use Chernoff’s bounds in the following setting.
@@
    $\mathbb{P}\left[B(n,p)\leq np/2\right]\leq e^{-np/8}$,
@@
    $\mathbb{P}\left[B(n,p)\geq 2np\right]\leq e^{-np/3}$.
@@
Suppose we have binary random variables $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ such that $\mathbb{P}\left[\mathbf{X}_{i}=1\right]\geq p$ regardless of the realisation of $\mathbf{X}_{1},\dots,\mathbf{X}_{i-1}$, for all $i\in[n]$. These variables are not necessarily independent, but we expect $\sum\mathbf{X}_{i}$ to be at least as large as a binomial variable of parameters $n,p$. We formalise this idea in Lemma 15.
@@
Let =$\mathcal{B}=\{B_{1},\ldots,B_{n}\}$ be a finite set of random (bad) events, all of which happen with probability at most $p$. Suppose that, for every $i\in[n]$, $B_{i}$ is independent from all but at most $d$ other events.
@@
In some applications, Lemma 16 is not sufficient: for these, we shall use the following lopsided version of the Symmetric Lovász Local Lemma (see e.g. [1]).
@@
Let =$\mathcal{B}=\{B_{1},\ldots,B_{n}\}$ be a finite set of random (bad) events, and let $d$ be a fixed integer. Suppose that, for every $i\in[n]$, there is a set $\Gamma(i)\subseteq[n]$ of size at most $d$ such that, for every $Z\subseteq[n]\setminus\Gamma(i)$,
@@ 3. A general upper bound for $\chi_{\rm pcf}^{h}$
Therefore, we perform a two-step colouring of the graph $G$. First, we shall use a modified version of the greedy algorithm presented in the proof of 3 to construct a proper precolouring $\sigma_{0}$ of $G$ such that all vertices of “small degree” have the required number of witnesses, while those of “large degree” may still miss one witness each. During the second step, we shall randomly recolour a fraction of the vertices with new colours in order to create one extra witness for each vertex of large degree. For that, we use a Rödl Nibble approach that takes inspiration from the work of Kamyczura and Przybyło [9], but incorporates new ideas to ensure that witnesses from the precolouring $\sigma_{0}$ are preserved, and also relies on a different procedure which we believe is conceptually simpler to understand and analyse.

### v2

@@ Abstract.
A proper $k$-colouring of a graph $G$ is called $h$-conflict-free if every vertex $v$ has at least $\min\,\{h,\deg(v)\}$ colours appearing exactly once in its neighbourhood. Let $\chi_{\rm pcf}^{h}(G)$ denote the minimum $k$ such that such a colouring exists. We show that for every fixed $h\geq 1$, every graph $G$ of maximum degree $\Delta$ satisfies $\chi_{\rm{pcf}}^{h}(G)\leq h\Delta+\mathcal{O}(h\log\Delta)$. This expands on the work of Cho et al., and improves a recent result of Liu and Reed in the case $h=1$. We conjecture that for every $h\geq 1$ and every graph $G$ of maximum degree $\Delta$ sufficiently large, the bound $\chi_{\rm pcf}^{h}(G)\leq h\Delta+1$ should hold, which would be tight. When the minimum degree $\delta$ of $G$ is sufficiently large, namely $\delta\geq\max\{100h,2000\log\Delta\}$, we show that this upper bound can be further reduced to $\chi_{\rm{pcf}}^{h}(G)\leq\Delta+\mathcal{O}(\sqrt{h\Delta})$. This improves a recent bound from Kamyczura and Przybyło when $\delta\leq\sqrt{h\Delta}$.
@@ 1. Introduction
For a positive integer $k$, let us denote $k]\coloneqq\{1,\dots,k\}$ the set of the first $k$ positive integers. A $k$-colouring of a graph $G$ is an assignment $\sigma\colon V(G)\rightarrow[k]$. A colouring $\sigma$ of $G$ is proper if $\sigma(u)\neq\sigma(v)$ for every edge $uv\in E(G)$. Given a colouring $\sigma$ and a vertex $v\in V(G)$, a witness of $v$ is a neighbour $u\in N(v)$ such that the colour $\sigma(u)$ appears exactly once in $N(v)$; we say that $\sigma(u)$ is a solitary colour of $v$ in $\sigma$. The colouring $\sigma$ is proper conflict-free (or simply pcf) if it is proper and every non-isolated vertex $v$ has a witness. We let $\chi_{\rm pcf}(G)$ be the smallest integer $k$ such that a pcf $k$-colouring of $G$ exists. The notion of pcf colouring was introduced by Fabrici et al. [8], and was first investigated in the class of graphs of maximum degree $\Delta$ by Caro, Petruševski and Škrekovski [3]. Among other results, they obtained the upper bound $\chi_{\rm pcf}(G)\leq 5\Delta(G)/2$, and conjectured that the optimal upper bound should actually be $\Delta(G)+1$.
@@
The condition $\Delta\geq 3$ is needed because the conjectured bound does not hold for cycles. Indeed $\chi_{\rm pcf}(C_{5})=5=\Delta(C_{5})+3$, and for every other cycle $C$ whose length is not a multiple of $3$, $\chi_{\rm pcf}(C)=4=\Delta(C)+2$. The condition that $G$ is connected ensures that there is no isolated vertex. The conjectured upper bound is best possible. For instance, if $G$ is a graph obtained by subdividing once any subset of edges from the complete graph $K_{n}$ with $n\geq 4$, then $\chi_{\rm pcf}(G)=n=\Delta(G)+1$, while $\chi(G)=2$. In particular, when $G$ is obtained by subdividing once every edge of $K_{n}$, the difference between $\chi(G)$ and $\chi_{\rm pcf}(G)$ is unbounded.
As a first step towards resolving ˜1, Cranston and Liu [6] observed that a simple greedy algorithm provides a pcf $(2\Delta+1)$-colouring of any graph $G$ of maximum degree $\Delta$. They moreoever proved that $\chi_{\rm pcf}(G)<1.656\Delta$ when $\Delta$ is large enough. More recently, Liu and Reed [12] proved that ˜1 holds asymptotically, by showing $\chi_{\rm pcf}(G)\leq\Delta+\mathcal{O}(\Delta^{2/3}\log\Delta)$ as $\Delta\to\infty$. Our first contribution is to improve this upper bound by eliminating the polynomial factor in the second-order term.
@@
Fix any linear ordering $(v_{1},\dots,v_{n})$ of $V(G)$. For a vertex $v\in V(G)$, let $\mathcal{W}(v)$ be the first $\min\,\{h,\deg(v)\}$ neighbours of $v$ in the ordering. For $i$ from $1$ to $n$, colour $v_{i}$ with the first colour which does not appear in $N(v_{i})$, nor in $\mathcal{W}(u)$ for $u\in N(v_{i})$. Let $\sigma$ be the colouring obtained at the end of this procedure. One easily checks that $\sigma$ is proper. Furthermore, for all $v\in V$, the vertices of $\mathcal{W}(v)$ are witnesses of $v$ (the fact that $\mathcal{W}(v)$ is coloured before $N(v)\setminus\mathcal{W}(v)$ is important), therefore $\sigma$ is $h$-pcf. Observe that the number of colour constraints when colouring $v_{i}$ is at most $(h+1)\deg(v_{i})$, therefore $\sigma$ uses at most $(h+1)\Delta+1$ colours. ∎
@@
Let $G$ be a graph of maximum degree $\Delta$, and let $1\leq h\leq\Delta-2$ be an integer. Then $\chi_{\rm pcf}^{h}(G)\leq(h+1)\Delta-1$.
For the specific case $h=\Delta-2$, they exhibit a graph $G$ of maximum degree $\Delta$ arbitrarily large such that $\chi_{\rm pcf}^{h}(G)=(h+1)(\Delta-1)$. Note that this last term is also equal to $h\Delta+1$. We consider the opposite end of the spectrum by considering $h$ a fixed constant, and make an asymptotic improvement on ˜4 in that case.
@@
$\chi_{\rm pcf}^{h}(G)\leq h\Delta+\mathcal{O}\left(h\log\Delta\right).$
˜2 is a direct corollary of ˜5 with $h=1$. The precise dependency hidden behind the term $\mathcal{O}(h\log\Delta)$ is explicitly stated in ˜24.
The notion of pcf colourings can also be seen as a strengthening of odd colourings, cf. [14, 4, 7]. For a graph $G$ and an integer $h\geq 1$, an $h$-odd colouring of $G$ is a proper colouring such that every vertex $v$ observes at least $\min\,\{h,\deg(v)\}$ colours in its neighbourhood that appear an odd number of times. Let $\chi_{o}^{h}(G)$ be the minimum number of colours required for an $h$-odd colouring of $G$, and $\chi_{o}(G)\coloneqq\chi_{o}^{1}(G)$. Clearly, $\chi_{o}^{h}(G)\leq\chi_{\rm pcf}^{h}(G)$ for every $h\geq 1$. A weaker version of ˜1 was first formulated in [2] for the odd chromatic number of graphs of maximum degree $\Delta$, and it was shown in [7, Theorem 3] that $\chi_{o}(G)\leq\Delta+\mathcal{O}\left(\log\Delta\right)$. In that sense, ˜2 is also a strengthening of this result.
If $h$ is a fixed integer and $G$ is a graph of maximum degree $\Delta$, then we have $\chi_{o}^{h}(G)\leq h\Delta+\mathcal{O}\left(h\log\Delta\right)$ as a direct consequence of ˜5. The first-order term of this upper bound was shown to be tight in [7, Proposition 28].
@@
In turn, ˜6 demonstrates that ˜5 is tight up to the second-order term, i.e. for every fixed integer $h\geq 1$, there exists a graph of maximum degree $\Delta$ arbitrarily large such that $\chi_{\rm pcf}^{h}(G)\geq h\Delta+1$. In light of these observations, we propose an extended version of ˜1.
@@
The lower bound provided by ˜6 relies on a graph construction with many vertices of small degree (namely, of degree $h+1$). When there is an additional restriction on the minimum degree of $G$, it is possible to derive smaller upper bounds on $\chi_{\rm pcf}^{h}(G)$. In particular, Liu and Reed [11] have recently proved that ˜1 holds for regular graphs in a strong sense. Their result can be stated as follows.
@@
The condition on the minimum degree in ˜8 is of course very restrictive. One could wonder what happens when this condition is relaxed. This has been considered by Kamyczura and Przybyło [10], who showed the following.
@@
Let $G$ be a graph of maximum degree $\Delta$ and minimum degree $\delta$. For every integer $h$ satisfying $20\log\Delta\leq h\leq\delta/75$, one has
@@
The condition on $h$ implies that $\delta>1500\log\Delta$. When $h<20\log\Delta$, an upper bound on $\chi_{\rm pcf}^{h}(G)$ can be obtained by relying on the monotonicity of $\chi_{\rm pcf}^{h}(G)$ in $h$. One has $\chi_{\rm pcf}^{h}(G)\leq\chi_{\rm pcf}^{\left\lceil 20\log\Delta\right\rceil}(G)$, and can obtain an upper bound on $\chi_{\rm pcf}^{h}(G)$ with an application of ˜9. A direct consequence of ˜9 is that, when $h$ is a fixed integer and $\delta\gg\log\Delta$, one has $\chi_{\rm pcf}^{h}(G)\leq\Delta+\mathcal{O}(\Delta\log\Delta/\delta)$. We improve this bound when the minimum degree is not too large.
@@
Let $G$ be a graph of maximum degree $\Delta$ and minimum degree $\delta$. For every integer $h$ satisfying $20\log\Delta\leq h\leq\delta/100$, one has
@@
The upper bound provided by ˜10 is smaller than that of ˜9 in the regime $\delta=\mathcal{O}(\sqrt{h\Delta})$, and performs best at $\delta=100h$. However, we suspect that there is still room for improvement in that regime. For matters of comparison, it was proved in [7, Corollary 6.3] that if $3\log\Delta\leq h\leq\delta/2$, then $\chi_{o}^{h}(G)\leq\Delta+\mathcal{O}(h)$, and we conjecture that a similar bound should hold for $h$-pcf colourings as well.
@@
1.1. Notations and outline
@@
The paper is organised as follows. In Section˜2, we provide probabilistic tools that we will rely on in our proofs. We consider general graphs in Section˜3 and prove ˜5. In Section˜4 we consider graphs of constrained minimum degree and prove ˜10.
@@ 2. Probabilistic tools
In our proofs, we shall mostly use Chernoff’s bounds in the following setting. Those respectively follow from Lemma˜12(i) and (ii) with $\delta\coloneqq 1/2$ and $\delta\coloneqq 1/4$.
@@
    $\mathbb{P}\left[B(n,p)\leq\frac{1}{2}np\right]\leq e^{-np/8}$,
@@
    $\mathbb{P}\left[B(n,p)\geq\frac{5}{4}np\right]\leq e^{-np/36}$.
@@
Suppose we have binary random variables $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ such that $\mathbb{P}\left[\mathbf{X}_{i}=1\right]\geq p$ regardless of the realisation of $\mathbf{X}_{1},\dots,\mathbf{X}_{i-1}$, for all $i\in[n]$. These variables are not necessarily independent, but we expect $\sum\mathbf{X}_{i}$ to be at least as large as a binomial variable of parameters $n,p$. We formalise this idea in Lemma˜15.
@@
Let $\mathcal{B}=\{B_{1},\ldots,B_{n}\}$ be a finite set of random (bad) events, each of which happens with probability at most $p$. Suppose that, for every $i\in[n]$, $B_{i}$ is independent from all but at most $d$ other events.
@@
In some applications, Lemma˜16 is not sufficient: for these, we shall use the following lopsided version of the Symmetric Lovász Local Lemma (see e.g. [1]).
@@
Let $\mathcal{B}=\{B_{1},\ldots,B_{n}\}$ be a finite set of random (bad) events, and let $d$ be a fixed integer. Suppose that, for every $i\in[n]$, there is a set $\Gamma(i)\subseteq[n]$ of size at most $d$ such that, for every $Z\subseteq[n]\setminus\Gamma(i)$,
@@
If $4pd\leq 1$, then $\mathbb{P}\left[\bigcap_{i\in[n]}\limits\overline{B_{i}}\right]>0$.
@@ 3. A general upper bound for $\chi_{\rm pcf}^{h}$
Therefore, we perform a two-step colouring of the graph $G$. First, we shall use a modified version of the greedy algorithm presented in the proof of ˜3 to construct a proper precolouring $\sigma_{0}$ of $G$ such that all vertices of “small degree” have the required number of witnesses, while those of “large degree” may still miss one witness each. During the second step, we shall randomly recolour a fraction of the vertices with new colours in order to create one extra witness for each vertex of large degree. For that, we use a Rödl Nibble approach that takes inspiration from the work of Kamyczura and Przybyło [10], but incorporates new ideas to ensure that witnesses from the precolouring $\sigma_{0}$ are preserved, and also relies on a different procedure which we believe is conceptually simpler to understand and analyse.